# Selecting samples of galaxies with fewer Fingers-of-God

### v2

@@ 6 Discussion and conclusions
In §4.2, we showed explicitly that cosmological constraints derived from fitting EFTofLSS models to redshift-space power spectrum multipoles are suceptible to bias when FoG are present. The reason for this was explained in §2, where we used a simplified model to argue that the clustering of central-satellite and satellite-satellite pairs is expected to escape perturbative control on scales where their power spectrum quadrupole is negative. In §4.1, we took this insight further: we simulated a typical galaxy sample by populating the Abacus $N$-body mocks with galaxies resembling BOSS CMASS LRGs using an HOD, and found that the satellite-satellite contribution crosses this threshold around $k\approx 0.1\,h\,\rm{Mpc}^{-1}$, whereas for central-satellite pairs this is closer to $k\approx 0.2\,h\,\rm{Mpc}^{-1}$. From this, we argued that, already by $k\approx 0.2\,h\,\rm{Mpc}^{-1}$, there are fully non-perturbative contributions to the power spectrum multipoles at the $\mathcal{O}(10\%)$ level per mode, growing rapidly with $k$. This suggests that beyond those scales, FoG should be considered a source of unmodelable systematic error that is better removed.

### v3

@@ 6 Discussion and conclusions
In §4.2, we showed explicitly that cosmological constraints derived from fitting EFTofLSS models to redshift-space power spectrum multipoles are suceptible to bias when FoG are present. The reason for this was explained in §2, where we used a simplified model to argue that the clustering of central-satellite and satellite-satellite pairs is expected to escape perturbative control on scales where their power spectrum quadrupole is negative. In §4.1, we took this insight further: we simulated a typical galaxy sample by populating the Abacus $N$-body mocks with galaxies resembling BOSS CMASS LRGs using an HOD, and found that the satellite-satellite contribution crosses this threshold around $k\approx 0.1\,h\,\rm{Mpc}^{-1}$, whereas for central-satellite pairs this is closer to $k\approx 0.2\,h\,\rm{Mpc}^{-1}$. From this, we argued that, already by $k\approx 0.2\,h\,\rm{Mpc}^{-1}$, there are fully non-perturbative contributions to the power spectrum multipoles at the $\mathcal{O}(10\%)$ level per mode, growing rapidly with $k$. This suggests that beyond those scales, FoG should be considered a source of unmodelable systematic error that is better removed [^20] Alternative models have recently been developed which pursue a non-perturbative treatment of the FoG (e.g. [76]). If these are shown to describe FoG in realistic ways, they will constitute an interesting complement to the removal we advocate for..
@@ Acknowledgments
This work was carried out on the territory of xučyun (Huichin), the ancestral and unceded land of the Chochenyo speaking Ohlone people, the successors of the sovereign Verona Band of Alameda County.